Two cars approach each other on A straight and level road. Car a is traveling at 75Km/h, due north and car B is traveling at 45Km/h, due soth. Both velocities are measured relative to the ground. What is the velocity of a car A relative to an observer in car B
1
Expert's answer
2020-02-24T10:48:16-0500

The velocity of a car A relative to an observer in car B:


"\bold{V=V_A-V_B}"

"\bold{V}=75-(-45)=120\frac{km}{h},\text{ due north}"